Andretti Company has a single product called a Dak. The company normally produces and sells 82,000 Daks each year at a selling price of $58 per unit. The company’s unit costs at this level of activity are given below:

Direct materials $ 8.50
Direct labor 10.00
Variable manufacturing overhead 2.00
Fixed manufacturing overhead 7.00 ($574,000 total)
Variable selling expenses 3.70
Fixed selling expenses 3.00 ($246,000 total)
Total cost per unit $ 34.20

A number of questions relating to the production and sale of Daks follow. Each question is independent.

4. Due to a strike in its supplier’s plant, Andretti Company is unable to purchase more material for the production of Daks. The strike is expected to last for two months. Andretti Company has enough material on hand to operate at 25% of normal levels for the two-month period. As an alternative, Andretti could close its plant down entirely for the two months. If the plant were closed, fixed manufacturing overhead costs would continue at 35% of their normal level during the two-month period and the fixed selling expenses would be reduced by 20% during the two-month period.

a. How much total contribution margin will Andretti forgo if it closes the plant for two months?

b. How much total fixed cost will the company avoid if it closes the plant for two months?

c. What is the financial advantage (disadvantage) of closing the plant for the two-month period?
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Andretti
Variable cost per unit
Direct Materials                   8.50 A
Direct Labor                10.00 B
Variable Manufacturing overhead                   2.00 C
Variable Selling Expenses                   3.70 D
Total Variable cost per unit                24.20 E=A+B+C+D
Sell Price Per unit                58.00 F
Contribution Per unit                33.80 G=F-E
Number of Units         82,000.00 H
Contribution amount 2,771,600.00 I=G*H
Fixed cost
Fixed manufacturing overhead      574,000.00 J
Fixed selling expenses      246,000.00 K
Total Fixed cost      820,000.00 L=J+K
Net Income 1,951,600.00 M=I-L
Ans 4 a Plant close
Number of units         13,666.67 V=H/12*2
Contribution Per unit                33.80 G
Contribution lost      461,933.33 W=U*G
Ans 4 b
Savings in Fixed manufacturing overhead by 65%       (62,183.33) X=J/12*2*65%
Savings in Fixed selling expenses by 20%         (8,200.00) Y=K/12*2*20%
Total fixed cost to be avoided      (70,383.33)
Net Financial disadvantage of closing the plant      391,550.00 Z=W+X+Y